the area of a rectangle patio A(p)=32 p-p^2 where A(p) is the area in square meters and p is the width of the rectangle in meters
a) what dimensions give the maximum area of the patio
b) what is the maximum area of the patio

Answer by htmentor(1343)

The area is maximized when dA/dp = 0 = 32 - 2p -> p = 16 m
Therefore the area is A(16) = 32*16 - 16^2 = 256 m^2
Since area = l*w, the length is also 16 m
Ans: length, width = 16 m; Area = 256 m^2